I need to solve these quadratic equations by factoring... p^2+8p+7=0, x^2=24-5x
m^2=4+3m,

Answer by ikdeep(189)

p^2+8p+7=0
(Here you must know how we make factors i.e. splitting of the middle term...if you might not know then contact me again and I will explain to you... )
p^2+(7+1)p+7=0
p^2 + 7p + 1p + 7= 0
taking common from first two i.e. p^2 + 7p and last two i.e. 1p + 7
variables...we get
p(p+7) + 1(p+7) = 0
(p+7)(p+1)............Answer
Try to solve other problem in the same way...
Still if you won't be able to solve further, contact me again.
hope this will help you
Please feel free to revert back for any further queries.